The length of a new tv screen is 14 in larger than the width and the diagonal is 26 inches. What are the length and width?
Answer by ewatrrr(24785)

w^2 + (w + 14)^2 = 26^2
2w^2 + 28w -480 = 0
w^2 + 14w - 240 = 0
(w -10)(w + 24) = 0
w = 10
L = 24